What strength fishing line is needed to stop a 19-lb salmon
swimming at 9.2 ft/s in a distance of 4.5 in.?

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Ahmed
 

Initial speed (vi) of the salmon is, vi = 9.2 ft/s

Final speed (vf) of the salmon is, vf = 0 ft/s

So the average speed of the salmon will be, vav = (vi + vf) /2

= (0 ft/s+9.2 ft/s)/2 =    = 4.6 ft/s

Time required (t) to stop the salmon is equal to the distance travelled (x) by the salmon divided by the average velocity (vav) of the salmon.

So, t = x/ vav

To obtain the time required (t) to stop the salmon, substitute 4.5 in for x and 4.6 ft/s for vav in the equation  t = x/ vav,

t = x/ vav

  =(4.5 in)/(4.6 ft/s) =(4.5 in×0.0833 ft/1 in)/(4.6 ft/s) = (0.38 ft) /(4.6 ft/s) = 8.3×10-2 s

The deceleration a of the salmon will be, a = Δvt.

To find out the   deceleration a of the salmon, substitute 9.2 ft/s for Δv and 8.3×10-2 s for Δt in the equation a = Δvt,

a = Δvt

  = (9.2 ft/s)/(8.3×10-2 s) =110 ft/s2

To find out the force on the salmon, substitute 19-lb for W, 110 ft/s2 for a, and 32 ft/s2 for g in the equation F = Wa/g,

F = Wa/g

= (19-lb) (110 ft/s2)/ (32 ft/s2) = 65.31 lb

Rounding off to two significant figures, the force will be 65-lb.

From the above observation we conclude that, the force required to stop the 19-lb salmon swimming at 9.2 ft/s in a distance of 4.5 in would be 65-lb.

 

Regards

Arun (askIITians forum expert)